1. Is sec (βˆ’πœ½) equal to sec 𝜽 or - sec 𝜽? How do you know?Β 
  2. Verify the identity 1 - sin2 x cot2 x = sin2 x. Is there more than one way to verify the identity? If so, tell which way you think is easier and why.Β 
  3. Describe what is wrong with the simplification shown.

cos x - cos x sinΒ² x = cos x - cos x (1 + cosΒ² x)

Β Β Β Β Β Β Β Β Β Β = cos x - cos x - cosΒ³ x

Β Β Β Β Β Β Β Β Β Β = -cosΒ³ x

4. John said π’”π’Šπ’ 𝒙 + 𝒄𝒐𝒔 𝒙 = 𝟐 has no solution. Do you agree with John? Explain why or why not?


Verify identity:

  1. π’„π’π’•Β²πœ½+𝟏 / π’„π’π’•Β²πœ½ ≑ π’”π’†π’„Β²πœ½Β 
  2. Β (π’„π’”π’„Β²πœ½ βˆ’ 𝟏)π’”π’Šπ’Β²πœ½β‰‘ π’„π’π’”Β²πœ½
  3. 𝟏 βˆ’ 𝒔𝒆𝒄 𝜢 𝒄𝒐𝒔 𝜢 ≑ 𝒕𝒂𝒏 𝜢 𝒄𝒐𝒕 𝜢 βˆ’ 1
  4. 𝒕𝒂𝒏 𝑨+𝒄𝒐𝒕 𝑨 / 𝒔𝒆𝒄 𝑨 𝒄𝒔𝒄 𝑨 ≑ 1
  5. 𝟏 + 𝟐 π’•π’‚π’Β²πœ½ ≑ π’”π’†π’„β΄πœ½ βˆ’ π’•π’‚π’β΄πœ½
1
Expert's answer
2021-03-31T13:53:17-0400

Solution.

1.

"\sec{(-\theta)}=\sec{\theta},"

because "\sec\theta=\frac{1}{\cos \theta}" and "\cos(-\theta)=\cos \theta."

2.

"1-\sin^2x\cot^2x=1-\sin^2x\frac{\cos^2x}{\sin^2x}=\newline
=1-\cos^2x=\sin^2x."

As for me, this way is easier, because it is a way of proving by definition.

3.

Β Β Β "\cos x-\cos x\sin^2x=\cos x-\cos x(1-\cos^2x)=\newline
=\cos x-\cos x+\cos^3x=\cos^3x."

4.

I agree with John, because this equation has no solution.

"\sin x+\cos x=2,"

"\sqrt{2}\sin(x+\frac{\pi}{4})=2,"

"\sin(x+\frac{\pi}{4})=\frac{2}{\sqrt{2}},"

and "\frac{2}{\sqrt{2}}>1."

This equation has no real solutions.

Verify identity:

1.

"\cot^2\theta+\frac{1}{\cot^2\theta}=\cot^2\theta+\tan^2\theta,\newline
\sec^2\theta=\frac{1}{\cos^2\theta}=1+\tan^2\theta."

But "1\neq \cot^2\theta." This equality is not an identity.

"\cot^2\theta+\frac{1}{\cot^2\theta}\neq\sec^2\theta."

2.

"(\csc^2\theta-1)\sin^2\theta=(\frac{1}{\sin^2\theta}-1)\sin^2\theta=\newline=\frac{1-\sin^2\theta}{\sin^2\theta}\sin^2\theta=1-\sin^2\theta=\cos^2\theta."

3.

"1-\sec\alpha\cos\alpha=0," and "\tan\alpha\cot\alpha-1=0."

So, "1-\sec\alpha\cos\alpha=\tan\alpha\cot\alpha-1."

4.

"\tan A+\frac{\cot A}{\sec A\csc A}=\tan A+\frac{\cos A\cdot \cos A\sin A}{\sin A}=
\newline=\tan A+\cos^2A=\frac{\sin A+\cos^3A}{\cos A}\neq 1."

This equality is not an identity.

5.

"\sec^4\theta-\tan^4\theta=(\sec^2\theta-\tan^2\theta)(\sec^2\theta+\tan^2\theta)=\newline
=\frac{1-sin^2\theta}{\cos^2\theta}\cdot\frac{1+sin^2\theta}{\cos^2\theta}=
\frac{\cos^2\theta+\sin^2\theta+\sin^2\theta}{\cos^2\theta}=1+2\tan^2\theta."